Show that {1, 2, 3, 4, . . .} and {2, 4, 6, 8, . . .} have the same cardinality.

Hint: find a mapping and show it is 1–1 and onto.



1
Expert's answer
2021-02-01T19:03:23-0500

Let "A=\{ 1,2,3,4..........\}" and "B=\{ 2,4,6,8......\}" .

Define a map,

"f:A\rightarrow B" by "f(x)=2x" .

Claim: "f" is one-one and onto .

Let "f(x)=f(y)"

"\implies 2x=2y"

"\implies x=y" ( Dividing both side by 2 )

Hence , "f" is one - one .

Again for each "y\in B" there exist a "\frac{y}{2}" in "A" ("\frac{y}{2}" exists because "B" is a set of even number )such that "f(\frac{y}{2})=2×\frac{y}{2}=y"

Hence "f" is onto .

Therefore "f" is one-one onto function.

Hence "A" and "B" have the same cardinality .